OpenStudy (anonymous):

Help Please!!!!!!!!!! Solve these equations using the addition property. 3x - 4y = 7 3x + 2y = -5 The lines intersect at ( __, __ )

OpenStudy (anonymous):

Um... I can help you

OpenStudy (anonymous):

You have to do it like this:\[(3x-4y=7)(-1)\] In order to find the y you have to multiply by a negative number. It becomes like this: -3x+4y=-7 3x+2y=-5 You sum and or subtract the numbers and you obtain: -3x+3x=0 4y+2y=6y -7-5=-12

OpenStudy (anonymous):

Then from here it's easy to see that y=-2. From this point onward you have to substitute y in one of the original equations in order to find the x

OpenStudy (anonymous):

Just like this: 3x-4y=7 3x-4(-2)=7 3x+8=7 3x=-1 x=-1/3

OpenStudy (anonymous):

So the answer to your question should be that the lines intersect at (-1/3,-2)

OpenStudy (anonymous):

Is it okay? Did you understand?